OpenStudy (samigupta8):

A bead rests in a smooth vertical ring of radius R Bead is given a slight push so that it slides down around the ring.At some instant bead is at angular position theta with vertical then choose correct option for the following 1).Acceleration of bead is horizontal at theta=π-arc cos1/3 2).Acceleration of bead is horizontal at theta=arc cos 1/3

OpenStudy (samigupta8):

@ganeshie8

OpenStudy (samigupta8):

Oops it was .. A bead rests on a smooth vertical fixed ring

OpenStudy (abhisar):

You may try posting it regular physics section for better response. Thanks.

OpenStudy (samigupta8):

Sir i have posted questions there!! And i were left with no choice so did it

OpenStudy (samigupta8):

@michele_laino

OpenStudy (samigupta8):

@Michele_Laino sir I came up with this equation 2cos^2theta+2costheta-1=0 And this gives us theta as arc cos(√3-1) And in the option I gave for the question none satisfy this...

OpenStudy (samigupta8):

@Michele_laino

OpenStudy (michele_laino):

is the initial position of the bead like this: 

OpenStudy (samigupta8):

Yeah!! Exactly

OpenStudy (michele_laino):

at generic position, we can write this:  the acceleration is: \[a = \left( { - \frac{{{v^2}}}{R}\cos \theta ,\quad \frac{{{v^2}}}{R}\sin \theta - g} \right)\]